# A Little Depth Goes a Long Way: The Expressive Power of Log-Depth Transformers

### v1

@@ Appendix C Graph Connectivity Proof
For $\ell\in\{1,\ldots,\lceil\log n\rceil\}$, layer $82\ell+$ 8 computes the predicate $B_{\ell}(i,j)$ at the first $n^{2}$ position $p=in+j$, as follows. It uses a head with query $\langle\phi(i),\phi(j)\rangle$. The keys in the second set of $n^{3}$ positions $q=n^{2}+i^{\prime}n^{2}+k^{\prime}n+j^{\prime}$ are set to $\langle\phi(i^{\prime}),\phi(j^{\prime})\rangle$ (recall that $\phi(i^{\prime})$ and $\phi(j^{\prime})$ are available in the residual stream at $q$) and the corresponding values are set to the boolean $C_{\ell}(i^{\prime},k^{\prime},j^{\prime})$, stored previously in the residual stream. The head thus attends uniformly to the $n$ padding positions that have coordinates $(i,k^{\prime},j)$ for various choices of $k^{\prime}$. It computes the average of their values, which equals $h=\frac{1}{n}\sum_{k^{\prime}=1}^{n}C_{\ell}(i,k^{\prime},j)$ as well as $1/(2n)$ using an additional head. We observe that $h\geq 1/n$ if there exists a $k^{\prime}$ such that $C_{\ell}(i,k^{\prime},j)=1$, and $h=0$ otherwise. These conditions correspond precisely to $B_{\ell}(i,j)$ being $1$ and 00, respectively. We compute $h-1/(2n)$ and store it in the residual stream. Similar to the proof of Lem. A.7, the feedforward layer reads $\sigma=\mathrm{sgn}(h-1/(2n))$, computes $z=(1+\mathrm{ReLU}(\sigma))/2$, and writes $z$ to the residual stream. The value $z$ is precisely the desired $B_{\ell}(i,j)$ as $\sigma$ is $1$ when $h\geq 1/n$ and 00 when $h=0$. As in Lem. A.7, the intermediate value $h-1/(2n)$ written to the residual stream can be recomputed and reset in the next layer. As before, the transformer replaces the value $B_{\ell-1}(i,j)$ stored previously in the residual stream with the newly computed value $B_{\ell}(i,j)$ by adding $\psi(B_{\ell}(i,j)-B_{\ell-1}(i,j))$ to the stream at the same coordinates.
@@
Appendix D Proofs for Width Scaling and Chain of Thought Claims
See 6.1
@@ Appendix D Proofs for Width Scaling and Chain of Thought Claims
By assumption, we can construct an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit family in which the transformer weights for sequence length $n$ are hardcoded as constants. Next, we can apply standard arguments (Merrill et al., 2022; Merrill & Sabharwal, 2023a, b) to show that the self-attention and feedforward sublayers can both be simulated by constant-depth threshold circuits, and the size remains polynomial (though a larger polynomial). Thus, any function computable by a constant-depth, polynomial-width transformer is in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. ∎
See 6.2
@@
The high-level idea is that a polynomial-size circuit can enumerate all possible $O(\log n)$-length chains of thought. Then, in parallel for each chain of thought, we construct a threshold circuit that simulates a transformer (Merrill & Sabharwal, 2023a) on the input concatenated with the chain of thought, outputting the transformer’s next token. We then select the chain of thought in which all simulated outputs match the correct next token and output its final answer. The overall circuit has constant depth, polynomial size, and can be shown to be $\mathsf{L}$-uniform. Thus, any function computable by a transformer with $O(\log n)$ chain of thought is in $\mathsf{TC}^{0}$. ∎
Appendix E Experimental Details

### v2

@@ Appendix D Graph Connectivity Proof
For $\ell\in\{1,\ldots,\lceil\log n\rceil\}$, layer $82\ell+$ 8 computes the predicate $B_{\ell}(i,j)$ at the first $n^{2}$ positions $p=in+j$, as follows. It uses a head with query $\langle\phi(i),\phi(j)\rangle$. The keys in the second set of $n^{3}$ positions $q=n^{2}+i^{\prime}n^{2}+k^{\prime}n+j^{\prime}$ are set to $\langle\phi(i^{\prime}),\phi(j^{\prime})\rangle$ (recall that $\phi(i^{\prime})$ and $\phi(j^{\prime})$ are available in the residual stream at $q$) and the corresponding values are set to the boolean $C_{\ell}(i^{\prime},k^{\prime},j^{\prime})$, stored previously in the residual stream. The head thus attends uniformly to the $n$ padding positions that have coordinates $(i,k^{\prime},j)$ for various choices of $k^{\prime}$. It computes the average of their values, which equals $h=\frac{1}{n}\sum_{k^{\prime}=1}^{n}C_{\ell}(i,k^{\prime},j)$ as well as $1/(2n)$ using an additional head. We observe that $h\geq 1/n$ if there exists a $k^{\prime}$ such that $C_{\ell}(i,k^{\prime},j)=1$, and $h=0$ otherwise. These conditions correspond precisely to $B_{\ell}(i,j)$ being $1$ and 00, respectively. We compute $h-1/(2n)$ and store it in the residual stream. Similar to the proof of Lemma 8, the feedforward layer reads $\sigma=\mathrm{sgn}(h-1/(2n))$, computes $z=(1+\mathrm{ReLU}(\sigma))/2$, and writes $z$ to the residual stream. The value $z$ is precisely the desired $B_{\ell}(i,j)$ as $\sigma$ is $1$ when $h\geq 1/n$ and 00 when $h=0$. As in Lemma 8, the intermediate value $h-1/(2n)$ written to the residual stream can be recomputed and reset in the next layer. As before, the transformer replaces the value $B_{\ell-1}(i,j)$ stored previously in the residual stream with the newly computed value $B_{\ell}(i,j)$ by adding $\psi(B_{\ell}(i,j)-B_{\ell-1}(i,j))$ to the stream at the same coordinates.
@@
Appendix E Proofs for Width Scaling and Chain of Thought Claims
See 3
@@ Appendix E Proofs for Width Scaling and Chain of Thought Claims
By assumption, we can construct an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit family in which the transformer weights for sequence length $n$ are hardcoded as constants. Next, we can apply standard arguments [Merrill et al., 2022, Merrill and Sabharwal, 2023a, b] to show that the self-attention and feedforward sublayers can both be simulated by constant-depth threshold circuits, and the size remains polynomial (though a larger polynomial). Thus, any function computable by a constant-depth, polynomial-width transformer is in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. ∎
See 4
@@
The high-level idea is that a polynomial-size circuit can enumerate all possible $O(\log n)$-length chains of thought. Then, in parallel for each chain of thought, we construct a threshold circuit that simulates a transformer [Merrill and Sabharwal, 2023a] on the input concatenated with the chain of thought, outputting the transformer’s next token. We then select the chain of thought in which all simulated outputs match the correct next token and output its final answer. The overall circuit has constant depth, polynomial size, and can be shown to be $\mathsf{L}$-uniform. Thus, any function computable by a transformer with $O(\log n)$ chain of thought is in $\mathsf{TC}^{0}$. ∎
Appendix F Experimental Details
@@
Compute.
Each training run was launched on a single GPU. We estimate that, together, these experiments took about 1000 GPU hours.
License.
The codebase of Merrill et al. [2024], which we used for data generation, has MIT license.